# All or None: Identifiable Linear Properties of Next-token Predictors in Language Modeling

### v1

@@ E.1 Paraphrases
Here, we analyze how paraphrastic aspects of textual data can be described with relational context-query-reply (${\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y$) strings for a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$. We start by providing a tentative definition of of praphrasatic sentences in terms of their entailed conditional log-probabilities for different pairs of next-tokens.
@@
We say that ${\bm{\mathrm{q}}}_{2}\in\mathsf{Seq}(\mathcal{A})$ on $\mathcal{Y}_{2}\subseteq\mathcal{A}$ is a paraphrase of ${\bm{\mathrm{q}}}_{1}\in\mathsf{Seq}(\mathcal{A})$ on $\mathcal{Y}_{1}\subseteq\mathcal{A}$ for the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ if there exists $\beta\neq 0$ such that, for all $y_{0},y_{1}\in\mathcal{Y}_{1}\subseteq\mathcal{A}$, there exists $\hat{y}_{0},\hat{y}_{1}\in\mathcal{Y}_{2}\subseteq\mathcal{A}$, for which it holds:
@@
$\log\exp\big{(}({\bm{\mathrm{g}}}(y_{1})-{\bm{\mathrm{g}}}(y_{0}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})\big{)}$ $=\beta\log\exp\big{(}({\bm{\mathrm{g}}}(\hat{y}_{1})-{\bm{\mathrm{g}}}(\hat{y}_{0}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})\big{)}$ (Substituting 1 for the conditional probabilities)
@@
have rank equal to $\mathrm{dim}(\Gamma_{1})$ and $\mathrm{dim}(\Gamma_{2})$, respectively. Then, we make use of these matrices with their transpose in 233, obtaining:
@@
where in the last line we used $({\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{G}}})}={\bm{\mathrm{P}}}_{\Gamma_{1}}$, since ${\bm{\mathrm{G}}}$ in 234 spans $\Gamma_{1}$, and we defined ${\bm{\mathrm{O}}}:=({\bm{\mathrm{G}}}^{\top})^{+}{\hat{\bm{\mathrm{G}}}}^{\top}$. Proceeding similarly but from $\hat{\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})={\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})$ we obtain a similar equation to 236 :
@@
have rank equal to $\mathrm{dim}(\mathcal{F}_{1})$ and to $\mathrm{dim}(\mathcal{F}_{2})$, respectively. Therefore,by substituting ${\bm{\mathrm{F}}}_{1}$ and ${\bm{\mathrm{F}}}_{2}$ in 236 and in 237, we obtain:
@@
We use this in 241 to obtain the following:
@@
This shows that $k_{1}\leq o\leq k_{1}$, meaning that $\mathrm{rank}({\bm{\mathrm{O}}})=\mathrm{dim}(\Gamma_{1})$. Similarly, taking the rank of 242 we obtain a similar implication for $k_{1}$ and $k_{2}$ (as obtained for $k_{2}$ and $k_{1}$ for 241), that is:

### v2

@@ E.1 Paraphrases
Here, we analyze how paraphrastic aspects of textual data can be described with relational context-query-reply (${\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y$) strings for a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$. We start by providing a tentative definition of paraphrastic sentences in terms of their entailed conditional log-probabilities for different pairs of next-tokens.
@@
We say that ${\bm{\mathrm{q}}}_{2}\in\mathsf{Seq}(\mathcal{A})$ on $\mathcal{Y}_{2}\subseteq\mathcal{A}$ is a paraphrase of ${\bm{\mathrm{q}}}_{1}\in\mathsf{Seq}(\mathcal{A})$ on $\mathcal{Y}_{1}\subseteq\mathcal{A}$ for the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ if (1) there exists $\beta\neq 0$ such that, for all $y_{0},y_{1}\in\mathcal{Y}_{1}\subseteq\mathcal{A}$, and (2) there exist $\hat{y}_{0},\hat{y}_{1}\in\mathcal{Y}_{2}\subseteq\mathcal{A}$, for which it holds:
@@
$\log\exp\big{(}({\bm{\mathrm{g}}}(y_{1})-{\bm{\mathrm{g}}}(y_{0}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})\big{)}$ $=\beta\log\exp\big{(}({\bm{\mathrm{g}}}(\hat{y}_{1})-{\bm{\mathrm{g}}}(\hat{y}_{0}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})\big{)}$ (Substituting Equation (1) for the conditional probabilities)
@@
have rank equal to $\mathrm{dim}(\Gamma_{1})$ and $\mathrm{dim}(\Gamma_{2})$, respectively. Then, we make use of these matrices with their transpose in Equation (237), obtaining:
@@
where in the last line we used $({\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{G}}})}={\bm{\mathrm{P}}}_{\Gamma_{1}}$, since ${\bm{\mathrm{G}}}$ in Equation (238) spans $\Gamma_{1}$, and we defined ${\bm{\mathrm{O}}}:=({\bm{\mathrm{G}}}^{\top})^{+}{\hat{\bm{\mathrm{G}}}}^{\top}$. Proceeding similarly but from $\hat{\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})={\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})$ we obtain a similar equation to Equation (240) :
@@
have rank equal to $\mathrm{dim}(\mathcal{F}_{1})$ and to $\mathrm{dim}(\mathcal{F}_{2})$, respectively. Therefore,by substituting ${\bm{\mathrm{F}}}_{1}$ and ${\bm{\mathrm{F}}}_{2}$ in Equation (240) and in Equation (241), we obtain:
@@
We use this in Equation (245) to obtain the following:
@@
This shows that $k_{1}\leq o\leq k_{1}$, meaning that $\mathrm{rank}({\bm{\mathrm{O}}})=\mathrm{dim}(\Gamma_{1})$. Similarly, taking the rank of Equation (246) we obtain a similar implication for $k_{1}$ and $k_{2}$ (as obtained for $k_{2}$ and $k_{1}$ for Equation (245)), that is: